the weight of a hollow body,made of gold,is 36.670g in air,while it is 33.865g in water. If the density of gold is 19.3 g/cc, find the volume of the hollow part of the body.(Density of water is 1 g/cc.

Grade:Upto college level

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear mikir

weight if body =36.67 g

weight of body in water =33.865 g

this reduction in weight is dut to boyancy force

Fb = 36.67 g -33.865 g

    = 2.805 g

but Fb=wait of water replace by body

        =1*V

 

V= 2.805cc

but this total volume

so volume of hollow part + volume of solid part =V=2.805 cc

for volume of solid part

  volume of solid part*density of gold=weight of body in air

       volume of solid part=36.670/19.3 

                                    =1.9 cc

 so volume of hollow part =2.805-1.9

                                    =.905 cc


Please feel free to post as many doubts on our discussion forum as you can. If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin